# Training Hybrid Deep Quantum Neural Networks for Efficient Reinforcement Learning

### v5

@@ Abstract
Quantum circuits embed data in a Hilbert space whose dimensionality grows exponentially with the number of qubits, allowing even shallow parameterised quantum circuits (PQCs) to represent highly–correlated probability distributions that are costly for classical networks to capture. Reinforcement-learning (RL) agents, which must reason over long-horizon, continuous-control tasks, stand to benefit from this expressive quantum feature space, but only if the quantum layers can be trained jointly with the surrounding deep-neural components. Current gradient-estimation techniques (e.g., parameter-shift rule) make such hybrid training impractical for realistic RL workloads, because every gradient step requires a prohibitive number of circuit evaluations and thus erodes the potential quantum advantage. We introduce qtDNN, a tangential surrogate that locally approximates a PQC with a small differentiable network trained on-the-fly from the same minibatch. Embedding qtDNN inside the computation graph yields scalable batch gradients while keeping the original quantum layer for inference. Building on qtDNN we design hDQNN-TD3, a hybrid deep quantum neural network for continuous-control reinforcement learning based on the TD3 architecture. On the high-dimensional Humanoid-v4 benchmark, our agent reaches a test return that surpasses classical TD3, SAC and PPO baselines trained with identical compute. To our knowledge this is the first PQC-enhanced policy that matches or exceeds state-of-the-art classical performance on Humanoid. qtDNN has the potential to reduce quantum-hardware calls significantly and is designed to be compatible with today’s NISQ devices. The method opens a path toward applying hybrid quantum models to large-scale RL and other gradient-intensive machine-learning tasks.
@@ 1 Introduction
Quantum computing has initially attracted significant attention for its potential to solve classically challenging problems by exploiting data representations in a Hilbert space that grows exponentially with the number of qubits, as discussed by Biamonte et al. (2017). Machine learning (ML) is an application field that could benefit greatly from quantum computing, as patterns are learned from increasingly large datasets (cf.  Vaswani et al. (2023); DeepSeek-AI et al. (2025b)) and more efficient methods are necessary to advance artificial intelligence (cf. Robison (2024); Samsi et al. (2023); Wells (2025)). Quantum machine learning (QML) aims at enhancing the efficiency and generalizability of learning beyond classical methods (see Table 1 in Biamonte et al. (2017) for known speedups of various subroutines). Although fault-tolerant quantum machines capable of running complex circuits without error are not yet available to exponentially accelerate computing with algorithms based on Harrow et al. (2009) or Brassard et al. (2000), the rapid development of noisy intermediate-scale quantum (NISQ) devices has inspired a new class of QML techniques designed to operate under noisy constraints (cf. Wang and Liu (2024)). On one hand, recent quantum supremacy experiments demonstrated that NISQ quantum devices can generate distributions where measured qubit states exhibit correlations that can not be simulated with reasonable classical compute (cf. Arute et al. (2019); Acharya et al. (2025); Gao et al. (2025)). On the other hand, a subfield of ML, reinforcement learning (RL) has become a central pillar of modern AI (cf. ACM (2025)), by estimating classical probability distributions that capture regularities in its interaction data (cf. Vaswani et al. (2023)), which allows an RL agent to refine its policy to attain high performance on target tasks (cf. Schulman et al. (2017); Haarnoja et al. (2018); DeepSeek-AI et al. (2025a)). Nontrivial highly correlated distributions seen in sampled output data (cf. Andersen et al. (2025)) from NISQ QPUs have the potential to better model complicated correlations in applications (cf. Lo et al. (2023); Huang et al. (2022); Vaswani et al. (2023)), allowing RL models to achieve higher rewards, data efficiency, and robustness. Recent work from Jin et al. (2025) has shown that SOTA quantum reinforcement learning (QRL) with PPO modified to incorporate parametrized quantum circuits (PQCs), could have similar performance as the classical baseline models from Raffin (2020) on simple reinforcement learning benchmarks in Towers et al. (2024). This hybrid quantum machine learning is realized as a hybrid deep quantum neural network (hDQNN). In this setup, a classical deep neural network first maps the input into a vector of real numbers, those numbers become the control parameters fed to the downstream PQC during the forward pass. The input data to the PQC is encoded into the Hilbert space spanned by the $N$ qubits, $\mathcal{H}_{N\text{q}}\cong\mathbb{C}^{2^{N}}$. Low-dimensional classical input and projected (sampled) output act as bottleneck layers. QRL has shown a lot of promise when it comes to tackling tricky reinforcement learning tasks, especially those involving high-dimensional state and action (discrete and continuous) spaces. (cf. Dong et al. (2008); Paparo et al. (2014); Dunjko et al. (2017); Jerbi et al. (2021); Wu et al. (2025)) However, a significant barrier for applying hDQNNs to such high-dimensional problems is the cost of back-propagation through the PQC to update weights of classical DNNs before it.
Although methods exist for estimating gradients of a given PQC (cf. Wang and Liu (2024)), these can be impractical for real quantum hardware when the input and output dimensions are large as they require a large number of sequential quantum circuit measurements and are further incompatible with mini-batch updates due to the lack of parallel processing in physical QPUs. The lack of high-throughput training strategies compatible with NISQ hardware limits QRL’s practical performance and applicability to real-world tasks.
In this work, we introduce hDQNN-TD3, a hybrid quantum neural network architecture that approaches this bottleneck using a novel approach: a classical surrogate that we named qtDNN will approximate the PQC’s gradient updates on the backward pass. We demonstrate the hDQNN-TD3 model on the Humanoid-v4 reinforcement learning benchmark with better performance over the widely adopted classical SOTA counterparts. This paper is organized as follows: we will introduce the method that enables scalable training of hDQNN with classical DNNs around PQCs in Sec. 2. Following this, Sec. 3 will introduce hDQNN-TD3 for continuous variable RL tasks and show the experimental results from Humanoid-v4 demonstrating that hDQNN-TD3 outperforms comparable classical methods as well as publicly reported models using SOTA RL architectures from Zhao (2024, 2023a, 2023b). The last Sec. 4 will discuss outlooks to train significantly larger models.
@@ 2.1 Back-propagation and the quantum bottleneck
Modern machine learning methods are established with highly efficient backpropagation-based training: each operation in the computation graph supplies analytical Jacobians, so the loss can be differentiated w.r.t. millions of parameters and "backpropagated" through them with a single reverse pass. The updated parameters are used for a new forward pass, which creates a new differentiable loss. This process repeats until the loss reaches a set convergence condition.
On the other hand, hDQNNs are typically developed using PQCs to act as functional layers (QL) without learnable parameters. The described flow is typically interrupted by a QL  [^1] QL: produces output from the expectation or most probable outcome over $S$ shots of PQC, parametrized by its input, executed on simulator or physical QPU.. Fig. 1a shows the typical layout: a classical neural network PreDNN produces PQC control parameters $\mathbf{q}_{\text{i}}$, a QL outputting a bitstring $\mathbf{q}_{\text{o}}$ after executing PQC $S$ times, and a second classical neural network PostDNN maps $\mathbf{q}_{\text{o}}$ to the final prediction. $d_{\text{c-link}}$ classical links directly connecting PreDNN and PostDNN are added on top of the QL. These direct links allow classical communication between the DNNs around the QL, to help the PostDNN properly handle the QL’s sampled output.
It is however challenging to backpropagate loss efficiently through QLs. Unlike Gaussian re-parameterization layers, a PQC (and QL) has no closed-form gradient. Various strategies have been developed in recent years (cf.  Wang and Liu (2024)). Most methods obtain the gradient vector of a PQC with respect to its input parameters by varying each input parameter and using repeated PQC executions to estimate its partial derivative with respect to each varied input. Such methods become quickly intractable as it must run $\mathcal{O}(I)$ extra circuits per output dimension, driving the hardware cost to $I\times O\times S\times N_{b}$ evaluations per mini-batch for the QL with $I$ and $Q$ input and output dimensions respectively (cf. Sec. A.1.2). Despite requiring many repeated measurements that scale with the size of input and output, methods based on direct PQC gradient estimation make it intrinsically unfeasible to batch-backpropagate in parallel, prohibiting efficient mini-batched training for a hDQNN, especially when it is applied to complex reinforcement learning.
@@
Figure 1: a, a typical hDQNN contains a PreDNN connected to a PostDNN via both $d_{\text{c-link}}$ direct connections and a quantum layer realized as a parametrized quantum circuit. The hDQNN inferences through the training dataset and records $(\mathbf{x},\mathbf{y}_{\text{pred}},\mathbf{y},\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})$ of each forward pass into a buffer $\mathcal{M}$. In an update step, a batch, $\mathcal{B}$,of $N_{\text{b}}$ entries is sampled from $\mathcal{M}$ for updating hDQNN parameters. b, $\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}$ from $\mathcal{M}$ is stored into a qtDNN Replay Buffer. $N_{\text{qt}}$ tiny-batches, $\{\mathcal{B}_{\text{qt}}\}$, are sampled from $\mathcal{B}$ are then used to update qtDNN towards approximating PQC in $\mathcal{B}$. c, the updated qtDNN is then used in the surrogate model $Q_{\text{qt}}$ in this update step to facilitate the batched backpropagation of loss that are used to update the PreDNN and PostDNN with fixed qtDNN.
We resolve this bottleneck by introducing the qtDNN, a quantum tangential space learning architecture that approximates the local behaviors of a QL with a classical DNN. The qtDNN is a small, differentiable surrogate network that locally mimics the PQC inside each mini-batch. It serves as a component of the hDQNN that substitutes the PQC whenever differentiability is needed.
The qtDNN is used during the training loop to provide a differentiable first-order approximation of PQC’s behaviors, local to each minibatch during the backpropagation. Meanwhile, the forward inference pass goes through the QLs. To make training the qtDNN more efficient, we add a replay buffer (qtDNN-buffer) to each QL that stores the input, $\mathbf{q}_{\text{i}}$, and sampled output, $\mathbf{q}_{\text{o}}$, of the quantum layer as a pair, $(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})$, from each minibatch. This qtDNN-buffer is then used to train the qtDNN repeatedly with $N_{\text{qt}}$ tiny-batches randomly sampled from this qtDNN-buffer. This accelerates qtDNN’s convergence to a moving approximation to its target quantum layer (cf. Haarnoja et al. (2018)).
@@
    Surrogate fit: From the sampled mini-batch =$\mathcal{B}=\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}$ we draw $N_{\mathrm{qt}}$ tiny-batches to minimize a binary-cross-entropy between $\text{qtDNN}(\mathbf{q}_{\text{i}})$ and the observed bitstrings $\mathbf{q}_{\text{o}}$.
@@ 2.3 Local gradient fidelity of qtDNN
Intuitively, the PQC is a smooth function of its parameters over the small region explored by one mini-batch. A feed-forward MLP can capture its first-order behavior. Theorem 1 states this precisely:
@@
Let $Q:\mathbb{R}^{d}\to\mathbb{R}^{k}$ be the deterministic function realised by a PQC under the expectation over $S$ shots, and suppose $Q$ is continuously differentiable on an open set $\mathcal{U}\subset\mathbb{R}^{d}$. Fix a mini-batch =$\mathcal{B}=\{x_{i}\}_{i=1}^{N_{b}}\subset\mathcal{U}$ lying in a closed ball of radius $r$ centered at $\bar{x}=\frac{1}{N_{b}}\sum_{i}x_{i}$. For every pair of tolerances $(\varepsilon_{1},\varepsilon_{2})>0$ there exists a feed-forward MLP $Q_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ such that:
@@
The first inequality is simply the qtDNN learning to approximate the PQC’s outputs correctly, while the second one shows that it learns to approximate the gradient of the PQC had it been differentiable i.e. approximate its effect during the backpropagation.
Intuition behind the proof.
Because the mini-batch $\mathcal{B}$ lies inside a small radius-$r$ ball, $Q$ is well-approximated by its first-order Taylor expansion around the centre $\bar{x}$. Thus, we (i) approximate the first-order Taylor polynomial $P(x){:=}Q(\bar{x}){+}J_{Q}(\bar{x})(x{-}\bar{x})$ over $\mathcal{B}$ by an MLP $Q_{\theta}$, and (ii) use Lipschitz bounds on the second derivative of $Q$ to control the Taylor remainder and its gradient. Replacing the PQC by that MLP therefore changes (i) forward activations by at most $\varepsilon_{1}$ and (ii) their Jacobian by at most $\varepsilon_{2}$, which in turn perturbs the gradient flowing into earlier layers by the same order of magnitude. See Appendix A.2 for full details.
@@ 3.1 Problem set-up and motivation
We now instantiate the qtDNN surrogate inside a continuous–control agent (Fig. 2a). The target task is Humanoid-v4 $(d_{\text{obs}}=376,d_{\text{act}}=17$), a long–horizon MuJoCo benchmark considered challenging for both classical and quantum methods ( Towers et al. (2024)). Classical baselines trained with identical compute on public HuggingFace repositories obtain final episode returns PPO $\approx 685$, SAC 5 742absent5742$\approx$ 5\,742≈ 5 742, TD3 5 306absent5306$\approx$ 5\,306≈ 5 306Zhao (2024),  Zhao (2023a),  Zhao (2023b)). We also saw that the best seed of our hybrid agent surpasses all three with 6 01160116\,0116 011 (see §3.5.1).
Figure 2: a, the typical exploration flow for the Agent to interact with the reinforcement learning objective’s target environment, $\mathcal{E}$, and obtain experience data to be stored in the replay buffer, $\mathcal{M}$. b, the particular modular model architecture used for the Actor Model, $A$, that allows the intermediate mapping $T$ that maps its input vector $\mathbf{q}_{\text{i}}$ to $\mathbf{q}_{\text{o}}$ to be implemented differently without changing the rest of the model design. This allows comparing learning performance of different quantum and classical implementations fairly.
Actor design. The actor consists of a classical pre-network PreDNN, a middle block $T$, and a post-network PostDNN (Fig. 2b). We instantiate $T$ as:
  1. 1.
    PQC (our hDQNN-TD3 model),
  2. 2.
    fully–connected (FC),
  3. 3.
    random bit generator (RBG), or
  4. 4.
    all-zero layer (0L)
to create a controlled ablation.
@@
For clarity we recap the four alternating phases that occur during training, the symbols follow the notation introduced in Subsec. 2.2.
@@
At every environment step the current state $\mathbf{s}\in\mathcal{S}$ is encoded by the PreDNN. Its output is split into (i) classical bypass features $d_{\text{c-link}}$, and (ii) the control vector $\mathbf{q}_{\mathrm{i}}$ that parameterises the PQC layer. Sampling the PQC yields the binary vector $\mathbf{q}_{\mathrm{o}}$, which the PostDNN transforms—together with the bypass features—into the continuous action $\mathbf{a}\in\mathcal{A}$. Applying $\mathbf{a}$ in $\mathcal{E}$ returns the next state $\mathbf{s}^{\prime}$, reward $r$, and a done flag. The transition tuple $(\mathbf{s},\mathbf{a},\mathbf{q}_{\mathrm{i}},\mathbf{q}_{\mathrm{o}},r,\mathbf{s}^{\prime})$ is appended to the replay buffer $\mathcal{M}$.
@@
Algorithm 1 exposes the training loop pseudo code for hDQNN-TD3.
Moreover, Fig. 6 in Appendix D.2 details the data flow inside one update epoch, making explicit which parameters are optimized at which stage.
Algorithm 1 hDQNN-TD3: TD3 actor with a PQC and qtDNN surrogate
1:Environment $\mathcal{E}$, replay buffer $\mathcal{M}$, actor $A$ (PQC inside), twin critics $C_{1},C_{2}$, target networks $A^{\prime}$, $C_{1}^{\prime},C_{2}^{\prime}$, update periods $N_{\text{A}}$, surrogate epochs $N_{\text{qt}}$
2:for episode $=1,2,\dots$ do
3:     collect experience: interact using $\mathbf{a}_{t}=A(\mathbf{s}_{t})+\mathcal{N}_{t}$, store $(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\mathbf{s}_{t+1})$ in $\mathcal{M}$
4:     if time to update then
5:         Sample mini-batch $\mathcal{B}$ of size $N_{\text{b}}$ from $\mathcal{M}$
6:         (a) update qtDNN: train surrogate $A_{\text{qt}}$ for $N_{\text{qt}}$ tiny-batches $\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}\in\mathcal{B}$
8:         if global step mod $N_{\text{A}}=0$ then
11:         end if
12:     end if
13:end for

### v6

# Training Hybrid Deep Quantum Neural Network for Efficient Reinforcement Learning

@@ Abstract
Quantum circuits embed data in a Hilbert space whose dimensionality grows exponentially with the number of qubits, allowing even shallow parameterised quantum circuits (PQCs) to represent highly-correlated probability distributions that are costly for classical networks to capture. Reinforcement-learning (RL) agents, which must reason over long-horizon, continuous-control tasks, stand to benefit from this expressive quantum feature space, but only if the quantum layers can be trained jointly with the surrounding deep-neural components. Current gradient-estimation techniques (e.g., parameter-shift rule) make such hybrid training impractical for realistic RL workloads, because every gradient step requires a prohibitive number of circuit evaluations and thus erodes the potential quantum advantage. We introduce qtDNN, a tangential surrogate that locally approximates a PQC with a small differentiable network trained on-the-fly from the same minibatch. Embedding qtDNN inside the computation graph yields scalable batch gradients while keeping the original quantum layer for inference. Building on qtDNN we design hDQNN-TD3, a hybrid deep quantum neural network for continuous-control reinforcement learning based on the TD3 architecture, which matches or exceeds state-of-the-art classical performance on popular benchmarks. The method opens a path toward applying hybrid quantum models to large-scale RL and other gradient-intensive machine-learning tasks.
@@ 1 Introduction
Quantum computing has attracted significant attention for its potential to address classically hard problems by exploiting Hilbert spaces that grow exponentially with the number of qubits (Biamonte et al., 2017). Machine learning (ML) is a domain that could benefit greatly, as patterns are extracted from increasingly large datasets (Vaswani et al., 2023; DeepSeek-AI et al., 2025b) and demand efficient methods for advancing AI (Robison, 2024; Samsi et al., 2023; Wells, 2025). Quantum machine learning (QML) aims to enhance learning efficiency and generalization beyond classical methods (see Table 1 in Biamonte et al. (2017)). Although fault-tolerant machines to run algorithms such as Harrow et al. (2009) or Brassard et al. (2000) are not yet available, rapid progress on noisy intermediate-scale quantum (NISQ) devices has spurred QML methods that operate under noise constraints (Wang and Liu, 2024). On one side, quantum supremacy experiments (Arute et al., 2019; Acharya et al., 2025; Gao et al., 2025) show NISQ devices can generate distributions intractable for classical hardware. On the other, reinforcement learning (RL) has become central in AI (ACM, 2025), modeling regularities in interaction data (Vaswani et al., 2023) to refine policies and achieve strong task performance (Schulman et al., 2017; Haarnoja et al., 2018; DeepSeek-AI et al., 2025a). The highly correlated distributions sampled from NISQ devices (Andersen et al., 2025) could help model complex correlations in RL (Lo et al., 2023; Huang et al., 2022; Vaswani et al., 2023), potentially improving reward, efficiency, and robustness.
Recent work has begun exploring quantum reinforcement learning (QRL). For example, Jin et al. (2025) modified PPO with parameterised quantum circuits (PQCs) and achieved performance similar to classical baselines (Raffin, 2020) on simple benchmarks (Towers et al., 2024). Hybrid deep quantum neural networks (hDQNNs) combine classical deep neural networks that map observations into real parameters, fed into PQCs operating in Hilbert space $\mathcal{H}_{N\text{q}}\cong\mathbb{C}^{2^{N}}$. Sampled outputs act as bottleneck features. Such architectures show promise for high-dimensional state and action tasks (Dong et al., 2008; Paparo et al., 2014; Dunjko et al., 2017; Jerbi et al., 2021; Wu et al., 2025). Yet a major barrier is the prohibitive cost of backpropagating through PQCs to train the surrounding classical DNNs.
1.1 Related Work
Gradient estimation techniques for PQCs (Wang and Liu, 2024) exist but remain impractical on hardware with large input/output dimensions, as they require many sequential circuit evaluations and cannot exploit mini-batch parallelism. This limits throughput and makes hDQNN training inefficient on real NISQ devices. Previous QRL studies demonstrated feasibility on small tasks (Jin et al., 2025), but lacked scalable training strategies. Classical surrogates have been proposed to mimic PQCs, yet typically either replace inference (losing quantum benefit) or lack guarantees of gradient fidelity.
1.2 Contributions
In this work we introduce hDQNN-TD3, a hybrid quantum actor-critic agent addressing the PQC back-propagation bottleneck via a new surrogate, qtDNN, which:
  1. 1.
    Provides a mini-batch-local differentiable approximation of the PQC used only in the backward pass, while the forward path and inference remain quantum.
  2. 2.
    Enables scalable mini-batch training by supplying gradients without additional quantum evaluations, compatible with high-throughput GPU pipelines.
  3. 3.
    Offers theoretical support through a local gradient-fidelity guarantee, ensuring that surrogate back-propagation perturbs upstream gradients only within controlled tolerances.
  4. 4.
    Demonstrates empirical effectiveness on the Humanoid-v4 benchmark, outperforming widely adopted classical baselines (Zhao, 2024; 2023a; 2023b) and validating the nontrivial modeling capacity of PQCs compared with fully connected, random, or zero-layer alternatives.
@@ 2.1 Back-propagation and the quantum bottleneck
Modern machine learning methods are established with back-propagation-based training: each operation in the computation graph supplies analytical Jacobians, so the loss can be differentiated w.r.t. millions of parameters and "backpropagated" through them with a single reverse pass.
On the other hand, hDQNNs are typically developed using PQCs to act as functional quantum layers (QL) without learnable parameters. The described flow is typically interrupted by a QL  [^1] QL: for each input control vector $\mathbf{q}_{\mathrm{i}}$, run the PQC for $S$ shots, aggregate the $S$ measured bit-strings $\{\mathbf{b}^{(s)}\}_{s=1}^{S}$, and return the per‑qubit empirical marginal vector $\hat{\bm{p}}(\mathbf{q}_{\mathrm{i}})\in[0,1]^{N}$ with entries $\hat{p}_{\ell}=\frac{1}{S}\sum_{s=1}^{S}b^{(s)}_{\ell}$.. Fig. 1a shows the typical layout: a classical neural network PreDNN produces PQC control parameters $\mathbf{q}_{\text{i}}$, a QL outputting the empirical marginal vector $\mathbf{q}_{\text{o}}=\hat{\bm{p}}\in[0,1]^{N}$ after $S$ shots, and a second classical neural network PostDNN maps $\mathbf{q}_{\text{o}}$ to the final prediction. $d_{\text{c-link}}$ classical links directly connecting PreDNN and PostDNN are added on top of the QL. These direct links allow classical communication between the DNNs around the QL, to help the PostDNN properly handle the QL’s sampled output.
It is however challenging to backpropagate loss efficiently through QLs. Unlike Gaussian re-parameterisation layers, a PQC has no closed-form gradient. Various strategies have been developed in recent years (cf. Wang and Liu (2024)). Most methods obtain the gradient vector of a PQC with respect to its input parameters by varying each input parameter and using repeated PQC executions to estimate its partial derivative with respect to each varied input. Such methods become quickly intractable as it must run $\mathcal{O}(I)$ extra circuits per output dimension, driving the hardware cost to $I\times O\times S\times N_{b}$ evaluations per mini-batch for the QL with $I$ and $O$ input and output dimensions respectively (cf. App. A.1.2). Despite requiring many repeated measurements that scale with the size of input and output, methods based on direct PQC gradient estimation make it intrinsically unfeasible to batch-backpropagate in parallel, prohibiting efficient mini-batched training for a hDQNN, especially when it is applied to complex reinforcement learning.
@@
Figure 1: a, a typical hDQNN contains a PreDNN connected to a PostDNN via both $d_{\text{c-link}}$ direct connections and a quantum layer realised as a parametrised quantum circuit. The hDQNN inferences through the training dataset and records $(\mathbf{x},\mathbf{y}_{\text{pred}},\mathbf{y},\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})$ of each forward pass into a buffer $\mathcal{M}$. In an update step, a batch, $\mathcal{B}$,of $N_{\text{b}}$ entries is sampled from $\mathcal{M}$ for updating hDQNN parameters. b, $\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}$ from $\mathcal{M}$ is stored into a qtDNN Replay Buffer. $N_{\text{qt}}$ tiny-batches, $\{\mathcal{B}_{\text{qt}}\}$, are sampled from $\mathcal{B}$ and then used to update qtDNN towards approximating PQC in $\mathcal{B}$. c, the updated qtDNN is then used in the surrogate model $Q_{\text{qt}}$ in this update step to facilitate the batched back-propagation of loss that are used to update the PreDNN and PostDNN with fixed qtDNN.
We resolve this bottleneck by introducing the qtDNN, a quantum tangential space learning architecture that approximates the local behaviours of a QL with a classical DNN. The qtDNN is a small, differentiable surrogate network that locally mimics the PQC inside each mini-batch. It serves as a component of the hDQNN that substitutes the PQC whenever differentiability is needed.
The qtDNN is used during the training loop to provide a differentiable first-order approximation of PQC’s behaviours, local to each mini-batch during the back-propagation. Meanwhile, the forward inference pass goes through the QLs. To make training the qtDNN more efficient, we add a replay buffer (qtDNN-buffer) to each QL that stores the input, $\mathbf{q}_{\text{i}}$, and sampled output, $\mathbf{q}_{\text{o}}$, of the quantum layer as a pair, $(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})$, from each mini-batch. This qtDNN-buffer is then used to train the qtDNN repeatedly with $N_{\text{qt}}$ tiny-batches randomly sampled from this qtDNN-buffer. This accelerates qtDNN’s convergence to a moving approximation to its target quantum layer (cf. Haarnoja et al. (2018)).
@@
    Surrogate fit: From the sampled mini-batch $\mathcal{B}=\{(\mathbf{q}_{\text{i}},\mathbf{q}_{\text{o}})\}$ we draw $N_{\mathrm{qt}}$ tiny-batches to minimize a binary-cross-entropy between $\text{qtDNN}(\mathbf{q}_{\text{i}})$ and the observed bit-strings $\mathbf{q}_{\text{o}}$.
@@
Note that we refit the qtDNN at every actor update on the current mini-batch and freeze it only for the remainder of that update epoch; a justification and ablations are given in App. B.3. For complexity comparisons against parameter‑shift and state‑vector simulators, see App. A.1.2.
@@ 2.3 Local gradient fidelity of qtDNN
Intuitively, the PQC is a smooth function of its parameters over the small region explored by one mini-batch. A feed-forward MLP can capture its first-order behaviour. Theorem 1 states this precisely:
@@
Let $Q:\mathbb{R}^{d}\to[0,1]^{N}$ be the deterministic map that returns the per‑qubit SS‑shot marginal vector of the PQC, and suppose $Q\in C^{1}$ on an open set $\mathcal{U}\subset\mathbb{R}^{d}$. Fix a mini-batch $\mathcal{B}=\{x_{i}\}_{i=1}^{N_{b}}\subset\mathcal{U}$ lying in a closed ball of radius $r$ centered at $\bar{x}=\frac{1}{N_{b}}\sum_{i}x_{i}$. For every pair of tolerances $(\varepsilon_{1},\varepsilon_{2})>0$ there exists a feed-forward MLP $Q_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ such that:
@@
The first inequality is simply the qtDNN learning to approximate the PQC’s outputs correctly, while the second one shows that it learns to approximate the gradient of the PQC had it been differentiable i.e. approximate its effect during the back-propagation [^2] We remark that Theorem 1 makes no assumption on the depth, connectivity, or gate layout of the underlying PQC; it guarantees local fidelity whenever the qtDNN fits the mini-batch outputs to within a target BCE loss. Empirical results suggest that this holds even for moderately deep and entangled circuits. We leave further tests on QAOA-style or chaotic ansätze as future work.. See App. B.1 for full proof details and App. D.1 for a pseudo-code constructing $Q$.
Our hardware‑efficient ansatz uses only smooth rotations with fixed entanglers; the unitary and per‑qubit expectations are analytic in the controls, hence $Q\in C^{1}$ on the region explored by a mini‑batch (cf. App. B.2)
@@ 3.1 Problem set-up and motivation
We now instantiate the qtDNN surrogate inside a continuous–control agent (Fig. 2a). The target task is Humanoid-v4 $(d_{\text{obs}}=376,d_{\text{act}}=17$), a long–horizon MuJoCo benchmark considered challenging for both classical and quantum methods ( Towers et al. (2024)). Classical baselines trained with identical compute on public HuggingFace repositories obtain final episode returns PPO $\approx 685$, SAC $\approx 5\,742$, TD3 $\approx 5\,306$Zhao (2024),  Zhao (2023a),  Zhao (2023b)).
Figure 2: a, the typical exploration flow for the Agent to interact with the reinforcement learning objective’s target environment, $\mathcal{E}$, and obtain experience data to be stored in the replay buffer, $\mathcal{M}$. b, the particular modular model architecture used for the Actor Model, AA, that allows the intermediate mapping $T$ that maps its input vector $\mathbf{q}_{\text{i}}$ to $\mathbf{q}_{\text{o}}$ to be implemented differently without changing the rest of the model design. This allows comparing learning performance of different quantum and classical implementations fairly.
Actor design. The actor consists of a classical pre-network PreDNN, a middle block $T$, and a post-network PostDNN (Fig. 2b). To create a controlled ablation, we instantiate $T$ as the following separate blocks: a PQC (our hDQNN-TD3 model); a fully–connected layer (FC); a random bit generator (RBG); and an all-zero layer (0L).
Note. While $T$ is instantiated as the above for the ablation, the actor always contains the PreDNN and the PostDNN. For instance, in our experiments, the “FC model” has 5 layers: 2 layers from the PreDNN, 1 layer for $T$, and 2 layers from the PostDNN.
@@
For clarity we recap the four alternating phases that occur during training, the symbols follow the notation introduced in Subsec. 2.2. See App. D.2 for the pseudo-code of the training loop.
@@
At every environment step the current state $\mathbf{s}\in\mathcal{S}$ is encoded by the PreDNN. Its output is split into (i) classical bypass features $d_{\text{c-link}}$, and (ii) the control vector $\mathbf{q}_{\mathrm{i}}$ that parameterises the PQC layer. Sampling the PQC yields the binary vector $\mathbf{q}_{\mathrm{o}}$, which the PostDNN transforms, together with the bypass features, into the continuous action $\mathbf{a}\in\mathcal{A}$. Applying $\mathbf{a}$ in $\mathcal{E}$ returns the next state $\mathbf{s}^{\prime}$, reward $r$, and a done flag. The transition tuple $(\mathbf{s},\mathbf{a},\mathbf{q}_{\mathrm{i}},\mathbf{q}_{\mathrm{o}},r,\mathbf{s}^{\prime})$ is appended to the replay buffer $\mathcal{M}$.
@@
Moreover, Fig. 7 details the data flow inside one update epoch, making explicit which parameters are optimised at which stage.